A 40,000 kg jet plane flies at a speed around 550 mph. ignoring any gravitational forces, What is the net force on the jet plane?



Give 2 situation wherein static is applicable in our lives.
Give 2 situations wherein kinetic friction is applicable in our lives.
1
Expert's answer
2021-10-06T08:13:37-0400

The net force can be found by dividing the power of the engines by the speed:


"F_\text{net}=\frac Pv=\frac P{550·1.609·1000/3600}=\frac P{246}\text{ N}."

2 situation wherein static is applicable in our lives: when we put a plate with food on the edge of the table and it doesn't fall, when we leave our PC mouse and it doesn't slide away.

2 situations wherein kinetic friction is applicable in our lives: when we pull bags with fertilizers in our garden, when we slide a book across our table without the need of lifting it.